User avatar
Vegas
4 Jan 2022 12:22 pm
User avatar
Beevee's Owner/Giant Slayer
Beevee's Owner/Giant Slayer
20,956 posts

So I was faced with rolling the dice on a 20% chance of an adverse outcome from the jab - and that number is conservative, since the jab was passed in haste through the waiver of ordinary hurdles. Or I could buck the tiger's odds on a disease that had a 3 in 10k chance of ending me. 


Are there any obvious flaws in my math or logic, dear reader? Image
It looks like you are spot on, but there is a little more math that needs to be done for no other reason than to prove it mathematically.  I believe (though I may be wrong) that to find the real probabilities, there are 4 cases: 
  1. vaccine with adverse conditions without infection
  2. Vaccine with adverse conditions with infection
  3. vaccine without adverse conditions with infection
  4. Vaccine without adverse conditions without infection
Probabilities: 
  • Vaccine with adverse conditions = .2
  • Vaccine without adverse conditions = .8
  • Infected = 3/10000 = .03%
  • Not infected = 99.97%
Case 1) .2 x .9997 = 19.994%
Case 2) .2 x .0003 = .006%
Case 3) .8 x .03 = 2.4%
Case 4) .8 x .9997 = 79.98%

Case 4 has the highest probability followed by Case 1. However, these are only the probabilities with the vaccine. I would say that case 2 is the most applicable to your argument. Very small probability, so you are spot on. 

 
Blackvegatble's hypcorisy summed up in one post: [/size]
Blackvegetable » 7 minutes ago » wrote: ↑7 minutes ago
Very simple questions...

From which you are running...



Image
Updated 4 minutes ago
© 2012-2026 Liberal Forum

Search